# Robust Score-Based Quickest Change Detection

### v1

@@ VII Identification of the Least Favorable Distributions
$\quad\min_{R_{\infty}\in\mathcal{G}_{\infty},R_{1}\in\mathcal{G}_{1}}D_{F}(R_{1}||R_{\infty})$ $=\min_{R_{\infty}\in\mathcal{G}_{\infty},R_{1}\in\mathcal{G}_{1}}\,\mathbb{E}_{R_{1}}\biggl{\{}\|\nabla_{x}\log r_{1}(X)-\nabla_{x}\log r_{\infty}(X)\|_{2}^{2}\biggr{\}}$
$\leq\quad\mathbb{E}_{G_{\hat{\mu}_{1}}}\biggl{\{}\|\nabla_{x}\log g_{\hat{\mu}_{1}}(X)-\nabla_{x}\log g_{\hat{\mu}_{\infty}}(X)\|_{2}^{2}\biggr{\}}=D_{F}(G_{\hat{\mu}_{1}}||G_{\hat{\mu}_{\infty}}).$ (34)
@@
$\mathbb{D}_{F}(G_{\mu_{1}}\|G_{\mu_{\infty}})=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\bigg{\|}\nabla_{x}\log g_{\mu_{1}}(X)-\nabla_{x}\log g_{\mu_{\infty}}(X)\bigg{\|}^{2}_{2}\bigg{\}}$
$=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\bigg{\|}\sum_{i=1}^{m}\beta_{\infty}^{i}(X)(X-\mu_{1}^{i})-\sum_{i=1}^{n}\beta_{1}^{i}(X)(X-\mu_{\infty}^{i})\bigg{\|}^{2}_{V}\bigg{\}}$
$=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{\}},$ (35)
@@
$\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{\}}\geq\mathbb{E}_{G_{\hat{\mu}_{1}}}\bigg{\{}\|\hat{\mu}_{\infty}-\hat{\mu}_{1}\|^{2}_{V}\bigg{\}}=\mathbb{D}_{F}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}}).$ (37)
Remark VII.5.
@@ VIII Numerical Simulations
We will consider Gaussian mixture models following the setup of Theorem VII.2 and a Gauss-Bernoulli Restricted Boltzmann Machine.
@@ VIII-A Gaussian Mixture Model Numerical Simulation
By Theorem VII.2, we know that $\mathcal{G}_{\infty}$ and $\mathcal{G}_{1}$ are disjoint. Furthermore, by Theorem VII.4, we can say that $R_{\infty}^{A},R_{1}^{A}$ are the least favorable distributions over $\mathcal{G}_{\infty},\mathcal{G}_{1}$ as it can be shown that their parameters are the nearest in $\mathbb{M}_{\infty},\mathbb{M}_{1}$ under the $V$-norm.
Next, we demonstrate that Assumption VI.1 holds for any $P_{\infty}\in\mathcal{G}_{\infty}$. Consider arbitrary Gaussian distributions with a common covariance matrix $V$. If $C=\mathcal{N}(\mu_{C},V)$ and $D=\mathcal{N}(\mu_{D},V)$, then $\mathbb{D}_{F}(C\|D)=\|V^{-1}(\mu_{C}-\mu_{D})\|^{2}$. Let us further restrict ourselves to Gaussians whose means can be expressed as $s\mathbf{1}_{2}$, where $s\in\mathbb{R}$ and $\mathbf{1}_{2}\in\mathbb{R}^{2}$ is a column vector of ones, and for any such Gaussian distribution $C$, let $s_{C}$ be the scalar constant such that $C=\mathcal{N}(s_{C}\mathbf{1},V)$. Then, we can write:
$\mathbb{D}_{F}(C\|D)=(s_{C}-s_{D})^{2}\|V^{-1}\mathbf{1}_{2}\|^{2}\propto(s_{C}-s_{D})^{2}.$ (39)
@@
R-AA $R_{\infty}^{A}$ $R_{1}^{A}$ Algorithm 1 -0.101 0.103
R-AB $R_{\infty}^{A}$ $R_{1}^{B}$ Algorithm 1 -0.622 0.104
R-BA $R_{\infty}^{B}$ $R_{1}^{A}$ Algorithm 1 -0.103 0.313
R-BB $R_{\infty}^{B}$ $R_{1}^{B}$ Algorithm 1 -0.618 0.310
N $R_{\infty}^{A}$ $R_{1}^{A}$ Algorithm 2 0.232 1.17
@@ VIII-B Gauss-Bernoulli Restricted Boltzmann Machine
To learn the scores of $Q_{\infty},Q_{1}$, we train $f_{\infty}(\cdot),f_{1}(\cdot)$. During training, a set of $N=1,000$ particles $X_{1},...,X_{N}$ are initially sampled from an arbitrary distribution of the post-change set $\mathcal{G}_{1}$ using Gibbs sampling ([19]). The network is then trained to minimize the following loss function:
@@
$\beta_{\infty}^{j}$ 9.21e-5 9.99e-1
$\beta_{1}^{j}$ 9.99e-1 6.22e-5
@@
To proceed with Algorithm 1, we need a method of calculating the Hyvärinen scores $\mathcal{S}_{H}(\cdot,Q_{\infty}),\mathcal{S}_{H}(\cdot,Q_{1})$. These Hyvärinen scores are the sum of a gradient log-density, defined in Equations 47 and 48, along with a Laplacian term. To estimate the Laplacian term of a mixture of GBRBM distributions, we use Hutchinson’s Trick ([25, 22]):
@@
R-00 $G_{\infty}^{0}$ $G_{1}^{0}$ Algorithm 1 -2.58 0.341
R-01 $G_{\infty}^{0}$ $G_{1}^{1}$ Algorithm 1 -2.59 1.18
R-10 $G_{\infty}^{1}$ $G_{1}^{0}$ Algorithm 1 -0.339 0.342
R-11 $G_{\infty}^{1}$ $G_{1}^{1}$ Algorithm 1 -0.324 1.19
N $G_{\infty}^{1}$ $G_{1}^{0}$ Algorithm 2 1.98 5.69
@@ IX Conclusion
In this work, we proposed the RSCUSUM algorithm. We defined the least favorable distributions in the sense of Fisher divergence. Using asymptotic analysis, we also analyzed the delay and false alarms of RSCUSUM in the sense of Lorden’s and Pollak’s metrics. We provided both theoretical and algorithmic methods for computing the least favorable distributions for unnormalized models. Numerical simulations were provided to demonstrate the performance of our robust algorithm.
@@ Appendix A Selection of Appropriate Multiplier
Note that $h(0)=0$, and $h^{\prime}(0)=\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})<0$ by Assumption VI.1. Next, we prove that either 1) there exists $\rho^{\star}\in(0,\infty)$ such that $h(\rho^{\star})=0$, or 2) for all $\rho>0$ we have $h(\rho)<0$.
@@
In practice, it is possible to use $m$ past samples $\mathbf{X}_{[-m+1,0]}$ to estimate the value of $\rho$. In particular, $\rho$ can be chosen as the positive root of the function $\rho\to\tilde{h}(\rho)$ given by
$\tilde{h}(\rho)\overset{\Delta}{=}\frac{1}{m}\sum_{i=1}^{m}[\exp(z_{\rho}(X_{i-m}))]-1.$
By Lemma VI.6 and its related technical discussions, the above equation has a root greater than zero with a high probability if $m$ is sufficiently large. In the case that $\rho$ is not chosen properly, the algorithm remains implementable but optimal performance of detection delay is not guaranteed.
It is worth noting that although our core results hold for a pre-selected $\rho$ that satisfied the condition discussed in Lemma VI.6, the effect of choosing any other $\rho^{\prime}$ amounts to the scaling of all the increments of RSCUSUM by a constant factor of $\rho^{\prime}/\rho$. This means that all of these results still hold adjusted for this scale factor. The result of Theorem VI.7 can be modified to be written as
$\mathbb{E}_{\infty}[T_{\texttt{RSCUSUM}}]\geq\exp\left\{\frac{\rho\omega}{\max(\rho,\rho^{\prime})}\right\},$
for any $\rho^{\prime}>0$.
It is easy to see that this scaling will change the statement of Theorem VI.8 accordingly to
$\mathbb{E}_{1}[T_{\texttt{RSCUSUM}}]\sim\frac{\max(\rho,\rho^{\prime})}{\rho}\frac{\log\gamma}{\rho^{\prime}(\mathbb{D}_{\texttt{F}}(P_{1}\|Q_{\infty})-\mathbb{D}_{\texttt{F}}(P_{1}\|Q_{1}))},$
as $\gamma\to\infty$.
In order to have the strongest results, we must choose $\rho$ as close to $\rho^{*}$ as possible.
@@ Appendix B Proofs of Delay and False Alarm Theorems
The theoretical analysis for delay and false alarms is analogous to the analysis from [15]. We again note that the proofs are similar but not identical as the pre-change distribution $P_{\infty}$ may not be equal to the least favorable distribution $Q_{\infty}$.
@@
The following proof uses similar arguments to the proof of Theorem 4 of [15]. We first introduce a technical definition in order to apply Corollary 2.2 of [18] to the proof of Theorem VI.8.
Definition B.1.
A distribution $P$ on the Borel sets of (,)($\infty$,$\infty$)( - ∞ , ∞ ) is said to be arithmetic if and only if it concentrates on a set of points of the form ±$\pm nd$, where $d>0$ and $n=1,2,\ldots$.
Remark B.2.
Any probability measure that is absolutely continuous with respect to the Lebesgue measure is non-arithmetic.
@@
Under the mild regularity conditions given by [16], which we assume in Assumption V.3:
@@
where $z(X)^{+}=\max(z(X),0)$. Additionally, $Q_{1}$ must be non-arithmetic in order to have Hyvärinen scores well-defined. Hence, by [18] Corollary 2.2.,

### v3

@@ VII Identification of the Least Favorable Distributions
$\quad\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}D_{F}(R_{1}||R_{\infty})$ $=\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\,\mathbb{E}_{R_{1}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log r_{1}(X)-\nabla_{x}\log r_{\infty}(X)\|_{2}^{2}\biggr{\}}$
$\leq\quad\mathbb{E}_{G_{\hat{\mu}_{1}}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log g_{\hat{\mu}_{1}}(X)-\nabla_{x}\log g_{\hat{\mu}_{\infty}}(X)\|_{2}^{2}\biggr{\}}=D_{F}(G_{\hat{\mu}_{1}}||G_{\hat{\mu}_{\infty}}).$ (39)
@@
$\mathbb{D}_{F}(G_{\mu_{1}}\|G_{\mu_{\infty}})=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\nabla_{x}\log g_{\mu_{1}}(X)-\nabla_{x}\log g_{\mu_{\infty}}(X)\bigg{\|}^{2}_{2}\bigg{\}}$
$=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\sum_{i=1}^{m}\beta_{\infty}^{i}(X)(X-\mu_{1}^{i})-\sum_{i=1}^{n}\beta_{1}^{i}(X)(X-\mu_{\infty}^{i})\bigg{\|}^{2}_{V}\bigg{\}}$
$=\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{\}},$ (40)
@@
$\mathbb{E}_{G_{\mu_{1}}}\bigg{\{}\frac{1}{2}\bigg{\|}\sum_{i=1}^{n}\beta_{1}^{i}(X)\mu_{\infty}^{i}-\sum_{i=1}^{m}\beta_{\infty}^{i}(X)\mu_{1}^{i}\bigg{\|}^{2}_{V}\bigg{\}}\geq\mathbb{E}_{G_{\hat{\mu}_{1}}}\bigg{\{}\frac{1}{2}\|\hat{\mu}_{\infty}-\hat{\mu}_{1}\|^{2}_{V}\bigg{\}}=\mathbb{D}_{F}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}}).$ (42)
Remark VII.6.
@@ VIII Numerical Simulations
We will consider Gaussian mixture models following the setup of Theorem VII.3 and a Gauss-Bernoulli Restricted Boltzmann Machine.
@@ VIII-A Gaussian Mixture Model Numerical Simulation
By Theorem VII.3, we know that $\mathcal{G}_{\infty}$ and $\mathcal{G}_{1}$ are disjoint. Furthermore, by Theorem VII.5, we can say that $R_{\infty}^{A},R_{1}^{A}$ are the least favorable distributions over $\mathcal{G}_{\infty},\mathcal{G}_{1}$ as it can be shown that their parameters are the nearest in $\mathbb{M}_{\infty},\mathbb{M}_{1}$ under the $V$-norm.
Next, we demonstrate that Assumption VI.1 holds for any $P_{\infty}\in\mathcal{G}_{\infty}$. Consider arbitrary Gaussian distributions with a common covariance matrix $V$. If $C=\mathcal{N}(\mu_{C},V)$ and $D=\mathcal{N}(\mu_{D},V)$, then $\mathbb{D}_{F}(C\|D)=\frac{1}{2}\|V^{-1}(\mu_{C}-\mu_{D})\|^{2}$. Let us further restrict ourselves to Gaussians whose means can be expressed as $s\mathbf{1}_{2}$, where $s\in\mathbb{R}$ and $\mathbf{1}_{2}\in\mathbb{R}^{2}$ is a column vector of ones, and for any such Gaussian distribution $C$, let $s_{C}$ be the scalar constant such that $C=\mathcal{N}(s_{C}\mathbf{1},V)$. Then, we can write:
$\mathbb{D}_{F}(C\|D)=\frac{1}{2}(s_{C}-s_{D})^{2}\|V^{-1}\mathbf{1}_{2}\|^{2}\propto(s_{C}-s_{D})^{2}.$ (44)
@@
R-AA $R_{\infty}^{A}$ $R_{1}^{A}$ Algorithm 1 -0.0518 0.0495
R-AB $R_{\infty}^{A}$ $R_{1}^{B}$ Algorithm 1 -0.0511 0.155
R-BA $R_{\infty}^{B}$ $R_{1}^{A}$ Algorithm 1 -0.311 0.0519
R-BB $R_{\infty}^{B}$ $R_{1}^{B}$ Algorithm 1 -0.309 0.157
N $R_{\infty}^{A}$ $R_{1}^{A}$ Algorithm 2 0.124 0.584
@@ VIII-B Gauss-Bernoulli Restricted Boltzmann Machine
To learn the scores of $Q_{\infty},Q_{1}$, we train $f_{\infty}(\cdot),f_{1}(\cdot)$. During training, a set of $N=100,000$ particles $X_{1},...,X_{N}$ are initially sampled from an arbitrary distribution of the post-change set $\mathcal{G}_{1}$ using Gibbs sampling [19]. The network is then trained to minimize the following loss function:
@@
$\beta_{\infty}^{j}$ 2.47e-6 1.00
$\beta_{1}^{j}$ 1.00 1.13e-7
@@
To proceed with Algorithm 1, we need a method of calculating the Hyvärinen scores $\mathcal{S}_{H}(\cdot,Q_{\infty}),\mathcal{S}_{H}(\cdot,Q_{1})$. These Hyvärinen scores are the sum of a gradient log-density, defined in Equations 52 and 53, along with a Laplacian term. To estimate the Laplacian term of a mixture of GBRBM distributions, we use Hutchinson’s Trick ([25, 22]):
@@
R-00 $G_{\infty}^{0}$ $G_{1}^{0}$ Algorithm 1 -2.59 0.362
R-01 $G_{\infty}^{0}$ $G_{1}^{1}$ Algorithm 1 -2.56 1.19
R-10 $G_{\infty}^{1}$ $G_{1}^{0}$ Algorithm 1 -0.324 0.357
R-11 $G_{\infty}^{1}$ $G_{1}^{1}$ Algorithm 1 -0.337 1.18
N $G_{\infty}^{1}$ $G_{1}^{0}$ Algorithm 2 2.10 5.81
@@ IX Conclusion
In this work, we proposed the RSCUSUM algorithm, which can detect a change in the distribution of an online high-dimensional data stream in real time. While the SCUSUM algorithm developed in our previous work could accomplish this by knowing only the score function of the pre- and post-change distributions, the new RSCUSUM method can accomplish this with the additional challenge of imprecisely knowing the pre- and post-change distribution score functions.
The RSCUSUM algorithm is defined using a novel notion of least-favorable distributions obtained using the Fisher divergence between pre- and post-change distributions. We also analyzed the average detection delay and the mean time to a false alarm of the algorithm using novel analytical techniques. We provided both theoretical and algorithmic methods for computing the least favorable distributions for unnormalized models, and provided an analytical solution in the case where the pre- and post- change models are specific Gaussian Mixture Models. Numerical simulations were provided to demonstrate the performance and implementability of our robust algorithm.
@@ Appendix A Selection of Appropriate Multiplier
Note that $h(0)=0$, and $h^{\prime}(0)=\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})<0$ by Assumption VI.1.
Next, we prove that either 1) there exists $\rho^{\star}\in(0,\infty)$ such that $h(\rho^{\star})=0$, or 2) for all $\rho>0$ we have $h(\rho)<0$.
@@ Appendix B Proofs of Delay and False Alarm Theorems
The theoretical analysis for the mean time to a false alarm is analogous to the analysis from [15]. We note that the proofs are similar but not identical as the pre-change distribution $P_{\infty}$ may not be equal to the least favorable distribution $Q_{\infty}$.
@@
Under the mild regularity conditions given by [16], which we assume in Assumption IV.4:
@@
where $z(X)^{+}=\max(z(X),0)$. Hence, by Wald’s lemma [18]